# Yrast states of quantum droplets confined in a ring potential

### v1

@@ I Introduction
One of the novel aspects of quantum droplets is the presence of an attractive and of a repulsive nonlinear term in the generalized Gross-Piteavskii equation that dictates the order parameter. Actually, it is the balance between these two terms which makes quantum droplets self-bound. Although the existence of a droplet does not require the presence of any trapping potential, clearly confining it in some potential may give rise to interesting effects.
@@
In what follows below we thus examine the yrast state of a quantum droplet, in a purely one-dimensional trapping potential, considering the so-called “symmetric” problem. From the analysis that is presented in the next sections, we distinguish between three regimes, depending on the value of the density of the homogeneous state $n_{0}=N/L$, with $N$ being the atom number and $L=2\pi R$ the circumference of the ring of radius $R$. The three regimes are characterized by a “low”, a “medium” and a “high” value of $n_{0}$. In the two opposite limits of low/high density, it is the attractive/repulsive nonlinear term that is the dominant one. Reference Lila examined the same transition using exact diagonalization for a system of few atoms. In the intermediate regime, the two terms are comparable and – clearly – this is the most interesting of the three regimes.
@@ IV.1 Droplets of “low” density
In this limit, the droplet carries the angular momentum via center-of-mass excitation, for all values of $\ell$. As a result, the density distribution is not affected by the angular momentum, as seen in Fig. 3. The corresponding energy per particle, i.e., the dispersion relation $E(\ell)/N$ (bottom plot in Fig. 3), has the exact form due to the excitation of the center of mass,

### v2

@@ I Introduction
One of the novel aspects of quantum droplets is the presence of an attractive and of a repulsive nonlinear term in the generalized Gross-Pitaevskii equation that dictates the order parameter. Actually, it is the balance between these two terms which makes quantum droplets self-bound. Although the existence of a droplet does not require the presence of any trapping potential, clearly confining it in some potential may give rise to interesting effects.
@@
In what follows below we thus examine the yrast state of a quantum droplet, in a purely one-dimensional trapping potential, considering the so-called “symmetric” problem (which is analysed below). One may wonder why we do not focus on the more general, asymmetric, problem; actually, this problem was considered in Ref. Steffi . The answer to this question is two-fold. First of all, although very interesting – the “asymmetric” problem is different than the one we consider here. The main reason is that, at least for a substantial population imbalance, the droplet coexists with a uniform background, since the excess atoms in the majority component can not bind to it. Furthermore, the case of a zero (i.e., very small in an actual experiment) population imbalance is experimentally relevant.
From the analysis that is presented in the next sections, we distinguish between three regimes, depending on the value of the density of the homogeneous state $n_{0}=N/L$, with $N$ being the atom number and $L=2\pi R$ the circumference of the ring of radius $R$. The three regimes are characterized by a “low”, a “medium” and a “high” value of $n_{0}$. In the two opposite limits of low/high density, it is the attractive/repulsive nonlinear term that is the dominant one. Reference Lila examined the same transition using exact diagonalization for a system of few atoms. In the intermediate regime, the two terms are comparable and – clearly – this is the most interesting of the three regimes.
@@ IV.1 Droplets of “low” density
In this limit, the droplet carries the angular momentum via center-of-mass excitation, for all values of $\ell$. As a result, the density distribution is not affected by the angular momentum, as seen in Fig. 3. The corresponding energy per particle, i.e., the dispersion relation $E(\ell)/N$ (bottom plot in Fig. 3), has the following, exact, form due to the excitation of the center of mass,
@@ VI Summary and overview
Last but not least, in the present study we focused on the symmetric case. Clearly, it would be interesting to consider also the asymmetric problem. In this case we no longer have one order parameter and a single equation, but rather two order parameters and two coupled nonlinear equations. Furthermore, at least when the population imbalance is not very small, we still have the formation of a droplet, while the extra atoms of the majority component surround the droplet. As a result, the two problems are rather different. Various questions arise immediately, including how the angular momentum is carried, what is the form of the dispersion relation in this case, etc. Reference Steffi has investigated this problem and has derived various interesting results, which, however, differ from the ones presented in this study.